x/2.4 = 3/8
Solve for X
Please show work

Answer by stanbon(75887)

x/2.4 = 3/8
----
Multiply both sides by 2.4 to get:
x = 2.4*3/8
----
x = 0.9
==============
Cheers,
Stan H.
--------------